# A deep reinforcement learning agent trained for interval timing exhibits similarities to biological systems

### v1

@@ Discussion
The DRL agent learned to perform the duration production task and exhibited neural patterns and behaviors that are comparable to those observed in biological systems and biologically plausible models of timing. First, it converged to an internal state with oscillatory neural activations, a ubiquitous pattern in biological neurons. Second, its action network majorly relied on high amplitude oscillating neurons that coincidentally peaked at reward generating time steps, akin to the coincidence detection mechanism in the SBF model of timing. Third, the agent initially learned its time-keeping mechanism through embodied interactions but later internalized it, enabling it to perform the timing task with different environmental inputs. This process parallels the evolution of biological mechanisms such as circadian rhythms, where an organism builds an internal time representation through interactions with the environment and ultimately uses it as a buffer to adapt to environmental changes

### v2

@@ Discussion
The DRL agent learned to perform the duration production task and exhibited neural patterns and behaviors that are comparable to those observed in biological systems and biologically plausible models of timing. First, it converged to an internal state with oscillatory neural activations, a ubiquitous pattern in biological neurons. Second, the emergent time-keeping mechanism exhibited high-level functional similarities with the SBF model of timing. Third, the agent initially learned its time-keeping mechanism through embodied interactions but later internalized it, enabling it to perform the timing task with different environmental inputs. This process parallels the evolution of biological mechanisms such as circadian rhythms, where an organism builds an internal time representation through interactions with the environment and ultimately uses it as a buffer to adapt to environmental changes
@@
Appendix
(a) Target duration: 2 time steps
(b) Target duration: 3 time steps
(c) Target duration: 5 time steps
Figure 7: PCA of the 256 LSTM hidden state neurons across time for the time production task with different target durations
(a) Target duration: 2 time steps
(b) Target duration: 3 time steps
(c) Target duration: 5 time steps
Figure 8: FFT of the 256 LSTM hidden state neurons across time for the time production task with different target durations
(a) Target duration: 2 time steps
(b) Target duration: 3 time steps
(c) Target duration: 5 time steps
Figure 9: Raw activations of the 256 LSTM hidden state neurons across time for the time production task with different target durations